### v2

# Spiking Meets Attention: Efficient Remote Sensing Image Super-Resolution with Attention Spiking Neural Networks

Abstract
Spiking neural networks (SNNs) are emerging as a promising alternative to traditional artificial neural networks (ANNs), offering biological plausibility and energy efficiency. Despite these merits, SNNs are frequently hampered by limited capacity and insufficient representation power, yet remain underexplored in remote sensing super-resolution (SR) tasks. In this paper, we first observe that spiking signals exhibit drastic intensity variations across diverse textures, highlighting an active learning state of the neurons. This observation motivates us to apply SNNs for efficient SR of RSIs. Inspired by the success of attention mechanisms in representing salient information, we devise the spiking attention block (SAB), a concise yet effective component that optimizes membrane potentials through inferred attention weights, which, in turn, regulates spiking activity for superior feature representation. Our key contributions include: 1) we bridge the independent modulation between temporal and channel dimensions, facilitating joint feature correlation learning, and 2) we access the global self-similar patterns in large-scale remote sensing imagery to infer spatial attention weights, incorporating effective priors for realistic and faithful reconstruction. Building upon SAB, we proposed SpikeSR, which achieves state-of-the-art performance across various remote sensing benchmarks such as AID, DOTA, and DIOR, while maintaining high computational efficiency. The code of SpikeSR will be available upon paper acceptance.
Index Terms:
Image super-resolution, spiking neural network, attention mechanism, remote sensing.
Figure 1: The visualization of pixel intensity and neuron voltage in images under various degradation factors reveals important insights. The pixel intensity map illustrates that the high-frequency components of the image tend to be smooth, indicating a reduction in sharp details during progressive downsampling. Neuron intensity maps, derived from a LIF model, show that high-frequency details persist with drastic fluctuations, suggesting that the neurons remain in an active state.
I Introduction
High-resolution remote sensing images (RSIs) contain fine-grained object structures and textures, which are critical for accurate interpretation in downstream tasks [1, 2, 3]. However, limited by the intrinsic resolution of airborne sensors, RSI can merely capture partial spatial details, resulting in suboptimal scene representation and visual quality. Image super-resolution (SR) aims to alleviate this problem by reconstructing high-resolution (HR) images from low-resolution (LR) observations [4, 5]. Despite this, SR remains a challenging ill-posed issue, as a degraded input may correspond to multiple plausible outputs.
Early efforts rely on hand-crafted priors to tame the ill-posedness, e.g., nonlocal mean [6, 7] and gradient profile [8], but they are often trapped in limited performance and scalability. Recent advances in artificial neural networks (ANNs), e.g., CNNs and Transformers, have witnessed remarkable progress in SR with large-capacity models [9, 10, 11, 12, 13]. However, they often come with a trade-off of increased computational overhead and growing storage costs, making them less efficient in practical scenarios, particularly when reconstructing large-scale RSIs.
More recently, brain-inspired spiking neural networks (SNNs), as the third generation of neural networks, have emerged as a promising alternative for energy-efficient intelligence [14, 15, 16]. Different from ANNs that encode features as continuous values, SNN can emulate biological communication with discrete spiking signals and propagate them by neurons, thus enjoying lower power consumption. As depicted in Fig. 1, our experiments reveal a novel finding that spiking neurons maintain an active learning state across LR RSIs, even in severely damaged textures. Specifically, we observed that degraded RSIs exhibit smoothed pixel intensities and obscured sharp details, posing a significant challenge to characterize high-frequency representations. In contrast, spiking signals retain drastic responses and pronounced spike rates, highlighting that neurons remain in an active learning state. This naturally arises a question: Can SNNs leverage their inherent properties to handle image degradation for efficient yet high-quality RSI SR?
In fact, to effectively grasp complex and diverse spatial details in RSIs, the network must possess adequate capacity and representation power. Unfortunately, there are two critical challenges when adapting SNNs for SR tasks. Firstly, spiking activity in SNNs inevitably causes pixel-wise information loss, which hampered the representation capacity of SNNs, especially when the network deepens. This stems from the discrete nature of binary spiking signals, leading to undesirable spiking degradation problems [17, 18]. Secondly, SNNs remain constrained by suboptimal membrane potential dynamics, restricting effective exploration of global context during spiking communications. This necessitates a customized strategy to optimize membrane potentials, but is barely explored before.
Figure 2: FLOPs and PSNR performance comparison. The circle sizes represent the number of parameters. Our SpikeSR outperforms SOTA efficient SR methods with high efficiency. PSNR results are averaged on the AID, DOTA, and DIOR datasets.
To address these limitations, we propose SpikeSR, an SNN-based framework inspired by human visual attention mechanisms, which can actively represent image degradation and modulate synaptic weights to focus on salient regions, which, in turn, regulate the spiking activity for improved capacity and representation power. Specifically, SpikeSR employs a concise yet effective spiking attention block (SAB) to optimize feature emphasis through spiking response dynamics, which integrates three key innovations: 1) the combination of CNN and SNN layers to mitigate information loss induced by discrete spiking activity; 2) introducing hybrid dimension attention (HDA) to recalibrate spiking response across both temporal and channel dimensions, facilitating a joint feature correlations learning; 3) accessing global self-similarity patterns in RSIs to infer spatial attention weights, incorporating effective priors for realistic and faithful reconstruction. Compared to state-of-the-art (SOTA) ANN-based efficient SR models, our SpikeSR demonstrates lower model complexity and superior performance, as shown in Fig. 2.
Our contributions are summarized as follows:
  • We pioneer an attention spiking neural network for efficient SR of remote sensing imagery, providing a new perspective on developing efficient models in large-scale Earth observation scenarios.
  • We devise a concise yet effective spiking attention block, which mitigates the information loss and regulates membrane potentials of spiking activity for improved representation of SNNs.
  • Extensive experimental results on various remote sensing datasets demonstrate that our SpikeSR achieves competitive SR performance against state-of-the-art ANNs-based methods.
Figure 3: Overall network architecture of SpikeSR. The LR input is replicated along the temporal dimension and then processed through a convolution to extract shallow features. The core module of SpikeSR is spiking attention groups (SAGs), which employ spiking attention blocks (SABs) to capture deep spiking representations. Each SAB contains three main components: (1) Spiking convolution block (SCB), (2) Hybrid dimension attention (HDA), and (3) Deformable similarity attention (DSA). The fusion block (FB) aggregates the spatial-temporal sequences, and pixelshuffle is used to reconstruct the SR output.
II Related Work
Deep Networks for SR. Inspired by the pioneering SRCNN [19], CNN-based SR methods have achieved remarkable progress, dominating the field for years. They mainly elaborated on the network design to tame the ill-posedness, with notable advances in residual connections [20, 21] and attention mechanisms [22, 23]. For instance, to grasp valuable self-similarity priors, self-similar attention [24] and non-local sparse attention [25] have been developed. However, these methods often suffer from high computational complexity due to exhaustive non-local modeling, making them less efficient in large-scale RSIs.
Recently, transformer-based SR models have demonstrated impressive performance, benefiting from their ability to model long-range dependencies. IPT [26] first introduces Transformers in SR field, but requires massive parameters and laborious pre-training processes. SwinIR [27] effectively reduces the model size by partitioning the image into smaller windows when applying multi-head attention mechanisms, while maintaining favorable performance. Despite these advancements, advanced SR models are often trapped by rising computational overhead and growing storage costs, posing significant concerns in real-world applications, particularly in remote sensing scenarios.
Efficient SR Models. To reduce computational budget, CARN [28] utilizes grouped convolutions and a cascading mechanism to improve the residual architecture. IMDN [29] progressively distills useful information during feature extraction and applies network pruning to further decrease complexity. FMEN [30] optimizes residual modules to accelerate inference. In Transformer-based SR methods, SPIN [31] enhances long-range modeling by combining self-attention with pixel clustering, facilitating interactions between superpixels. HiT-SR [32] expands the self-attention receptive field by applying different window sizes of hierarchal layers. Despite these successes, there is still room to further boost SR performance. Moreover, the potential of energy-efficient SNNs for SR tasks remains largely unexplored.
Spiking Neural Networks. Recent advances in neuromorphic computing have shown the great potential of SNNs in computational efficiency and power as CNNs. Currently, SNNs have been successfully applied to various tasks, such as image classification [14, 33], object detection and tracking [34, 15], optical flow estimation [35, 16], etc.
A common solution to build SNNs is converting pre-trained ANN models [36]. Li et al. [37] proposed a layer-wise calibration to minimize activation mismatch during conversion. Ding et al. [38] replaced ReLU with the rate norm layer, enabling direct conversion from a trained ANN to an SNN. Stockl et al. [39] used time-varying multi-bit spikes to better approximate activation functions. However, conversion-based methods face accuracy gaps and high latency due to extensive time-step simulations, resulting in increased latency and energy consumption.
An alternative involves using agent gradient functions for continuous relaxation of non-smooth spike activities, enabling direct training via backpropagation through time. Lee et al. [40] treated membrane potential as a signal to overcome discontinuities, enabling direct training from spikes. Wang et al. [41] introduced an iterative LIF model and proposed spatiotemporal backpropagation (STBP) based on approximate peak activity derivatives. Later, Zheng et al. [42] proposed temporal delay batch normalization (tdBN), which significantly enhanced the depth of SNNs. To bridge the performance gap between ANNs and SNNs, some methods borrowed insights from CNNs, applying residual learning [18, 43] and attention mechanisms [44, 45] to SNNs. Nonetheless, there has been limited exploration of pixel-level regression tasks, such as SR.
Figure 4: The illustration of our DSA. Note that we set the diagonal elements of the similarity matrix to zero before selecting the indices of the highest scores. The deformable convolution operates at the patch level, alleviating the mismatch between the most similar patches.
III Proposed Method
The architecture of the proposed SpikeSR is illustrated in Fig. 3, which mainly consists of spiking attention groups (SAGs). Before SAGs, we utilize a $3\times 3$ convolution to extract high-dimensional features from the LR input. These features are then processed through $m$ stacked SAGs to explore deeper representations. Each SAG includes $n$ spiking attention blocks (SABs), a spiking convolution block (SCB), and a residual connection. In the SCB, leaky integrate-and-fire (LIF) neurons [46, 47] are used to convert the inputs into binary spike sequences (i.e., 0 or 1). As shown in Fig. 3, the output of the LIF neuron is 1 when the membrane potential exceeds the threshold, and 0 otherwise. To optimize the membrane potential, we introduce hybrid dimension attention (HDA), which refines the spiking activity using an efficient temporal-channel joint attention [48]. Furthermore, the proposed deformable similarity attention (DSA) is employed to introduce global context for accurate SR. After the terminal SAG, a fusion block (FB) is utilized to convert the spike sequence features into continuous values. Finally, SpikeSR generates super-resolved output from the fused features by applying pixel-shuffle [49] and a $3\times 3$ convolutional layer.
III-A Spiking Attention Block
As evidenced in Fig. 1, regions degraded by different factors exhibit noticeable fluctuations when encoded by LIF, highlighting pronounced firing spike rates of neurons. This provides robust and latent informative spiking cues from LR images. Unlike ANNs that encode images into continuous decimal values, SNNs use discrete binary spike values for neuronal communication, and thus demonstrated undesirable information loss [50, 44], resulting in limited capacity to represent degraded LR images. To address this, the design philosophy of the SAB is focused on leveraging CNNs and attention mechanisms to regulate membrane potentials, facilitating high-quality feature representation for SR, which in turn affects the spiking activity.
As shown in Fig. 3, in particular, the output of the $n$-th SAB at the $t$-th time step is denoted as $\mathbf{X}^{t,n}$, and can be obtained by the following:
${\mathbf{X}^{t,n}}={\mathbf{X}^{t,n-1}}+{\rm DSA}({\rm HDA}({\bf{\bar{X}}}_{1}^{t,n}+{\bf{\bar{X}}}_{2}^{t,n})),$ (1)
where $\mathbf{\bar{X}}_{1}^{t,n}$ and $\mathbf{\bar{X}}_{2}^{t,n}$ are two feature representations obtained from parallel branches, defined by:
$\begin{array}[]{l}\mathbf{\bar{X}}_{1}^{t,n}={\rm SCB}({\rm SCB}({\mathbf{X}^{t,n-1}})),\\ \mathbf{\bar{X}}_{2}^{t,n}={\rm tdBN}({\rm Conv}({\mathbf{X}^{t,n-1}})),\end{array}$ (2)
where $\rm Conv$ represents a $3\times 3$ convolution layer, and $\rm tdBN$ means the threshold-dependent batch normalization.
Different from previous works that focus solely on separate temporal and channel modulation [17, 44, 51], SAB enables joint temporal-channel adjustment of the spike response with HDA, effectively achieving interdependencies between the temporal and channel scopes.
III-B Deformable Similarity Attention
Non-local self-similarity has been recognized as an effective prior for SR tasks [52]. However, existing non-local attention mechanisms are computationally expensive due to exhaustive non-local operations, which impedes their efficiency in large-scale RSIs. In contrast, the proposed DSA efficiently grasps complex self-similar patterns in RSIs at the patch level to infer intricate spatial weights. Then, we utilize the cross-attention (CA) paradigm to enhance long-range communication, facilitating the fusion of useful context.
The details of DSA are shown in Fig. 4. Since the object scale exhibits explicit diversity in RSIs, the input feature $F$ is downsampled using bilinear interpolation, forming a multi-scale feature pyramid. For clarity, we demonstrate this process by dividing the initial features into 9 patches. The final DSA exploits a cascaded patch division strategy [53]. Specifically, each patch is first average pooled to capture its spatial characteristics, then reshaped and subjected to self-similarity computation, yielding a similarity matrix. The final self-similarity scores are fused via matrix multiplication to enhance the multi-scale representation. The best-matching patch $\mathbf{\bar{P}}_{i}$} with $\mathbf{P}_{i}$} can be obtained by:
${\mathbf{\bar{P}}_{i}}=\mathop{\rm argmax}\limits_{{\mathbf{P}_{j}}}E{({\mathbf{P}_{i}})^{\rm T}}E({\mathbf{P}_{j}}),~~j\neq i,$ (3)
where we adopt the Gumbel-Softmax [54] to achieve the non-differentiable $\rm argmax$ function, and $E$ means the operation of average pooling and feature reshaping.
Although matched patches contain highly relevant similarity, they are inevitably subject to mismatches and geometric transformations. Hence, we use deformable convolution (DConv) to reduce the generation of hallucinated textures. The deformable feature $\mathbf{F}^{\rm D}$ at location $p_{0}$ is computed as follows:
${{\mathop{\mathbf{F}}\nolimits}^{\mathop{\rm D}\nolimits}}({p_{0}})=\sum\limits_{{p_{m}}\in{\rm{{\cal R}}}}{\omega({p_{m}})\cdot{\mathop{\mathbf{F}}\nolimits}({p_{0}}+{p_{m}}+\Delta{p_{m}})},$ (4)
[…]
$\mathbf{\bar{V}}={\rm softmax}(\mathbf{Q}{\mathbf{K}^{\rm T}}/\sqrt{d})\mathbf{V},$ (5)
Finally, the fused features are summarized with the original features $\mathbf{F}$ and fed into the multilayer perceptron (MLP), which contains two fully-connected layers, to obtain the final output:
$\mathbf{\tilde{F}}={\rm MLP}(\mathbf{F}+\mathbf{\bar{V}}).$ (6)
TABLE I: Quantitative comparison of SpikeSR with state-of-the-art methods on three remote sensing datasets. FLOPs are measured corresponding to an LR image of $160\times 160$ pixels.
Methods #Param. FLOPs AID [55] DOTA [56] DIOR [57] Average
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
Bicubic - - 28.86 0.7382 31.16 0.7947 28.57 0.7432 29.53 0.7587
SRCNN [19] 20K 0.512G 29.70 0.7741 32.10 0.8264 29.49 0.7768 30.43 0.7924
VDSR [20] 667K 17.08G 30.44 0.8004 33.22 0.8569 30.36 0.8036 31.34 0.8203
EDSR [21] 1518K 50.77G 30.65 0.8086 33.64 0.8648 30.63 0.8116 31.64 0.8283
CARN [28] 1112K 40.39G 30.66 0.8068 33.66 0.8633 30.64 0.8102 31.65 0.8268
IMDN [29] 715K 18.18G 30.71 0.8076 33.70 0.8641 30.73 0.8115 31.71 0.8277
RFDN-L [58] 681K 16.49G 30.69 0.8074 33.73 0.8642 30.72 0.8114 31.71 0.8277
LatticeNet [59]  [^7] K 19.39G 30.73 0.8089 33.75 0.8653 30.75 0.8126 31.74 0.8289
HNCT [60] 364K 8.48G 30.79 0.8104 33.83 0.8664 30.80 0.8136 31.81 0.8301
FMEN [30] 1046K 26.72G 30.65 0.8063 33.66 0.8631 30.66 0.8104 31.66 0.8266
RLFN [61] 544K 13.25G 30.70 0.8074 33.69 0.8636 30.70 0.8110 31.70 0.8273
ESRT [62] 752K 26.06G 30.77 0.8102 33.75 0.8668 30.81 0.8142 31.78 0.8304
SwinIR-light [27] 897K 23.56G 30.83 0.8114 33.94 0.8677 30.85 0.8149 31.87 0.8313
Omni-SR [63] 2803K 70.98G 30.89 0.8142 33.94 0.8695 30.89 0.8170 31.91 0.8336
NGswin [64] 995K 12.73G 30.79 0.8107 33.87 0.8667 30.79 0.8140 31.82 0.8305
SPIN [31]  [^5] K 18.91G 30.78 0.8098 33.85 0.8673 30.82 0.8139 31.82 0.8303
HiT-SR [32] 792K 21.04G 30.87 0.8138 33.93 0.8689 30.89 0.8167 31.90 0.8331
SpikeSR-S (Ours) 472K 15.21G 30.86 0.8126 33.89 0.8687 30.89 0.8162 31.88 0.8325
SpikeSR-M (Ours) 763K 24.00G 30.88 0.8133 33.92 0.8689 30.90 0.8163 31.90 0.8328
SpikeSR (Ours) 1042K 33.05G 30.91 0.8142 33.98 0.8700 30.95 0.8175 31.95 0.8339
III-C Fusion Block
To transform discrete spiking sequences into continuous pixel values, a common approach is to apply mean sampling along the time dimension. However, this naive process may lead to the loss of crucial spatial details, potentially affecting the SR quality. Therefore, we introduce a fusion block that adaptively aggregates spiking sequences and mitigates information loss. Given an input spike input $\mathbf{Y}$, the computation process of FB can be formulated as:
$\begin{array}[]{l}\mathbf{Y}_{1}=\sigma({\rm TA}(\mathbf{Y}))\otimes\mathbf{Y},\\ \mathbf{Y}_{2}=\sigma({\rm SA}(\mathbf{Y}))\otimes(1-{\mathbf{Y}_{1}}),\end{array}$ (7)
where $\rm TA$ and $\rm SA$ denote temporal and spatial attention [44], $\sigma$ means a sigmoid function and $\otimes$ denotes feature multiplication. The final output of FB is obtained by summing $\mathbf{Y}_{1}$ and $\mathbf{Y}_{2}$.
TABLE II: Quantitative comparison of SpikeSR with state-of-the-art methods on 30 scene types of AID datasets.
Scene types EDSR [21] CARN [28] IMDN [29] ESRT [62] SwinIR [27] SPIN [31] HiT-SR [32] SpikeSR
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
Airport 29.93 0.8282 29.96 0.8264 30.00 0.8270 30.09 0.8292 30.14 0.8307 30.11 0.8295 30.21 0.8325 30.27 0.8336
Bare Land 36.94 0.8837 36.92 0.8829 36.93 0.8834 36.90 0.8840 36.99 0.8841 36.96 0.8843 37.00 0.8846 36.98 0.8846
Baseball Field 33.05 0.8765 33.06 0.8753 33.17 0.8763 33.21 0.8773 33.27 0.8782 33.14 0.8759 33.29 0.8791 33.32 0.8794
Beach 34.18 0.8727 34.27 0.8737 34.29 0.8739 34.35 0.8755 34.39 0.8756 34.36 0.8757 34.38 0.8762 34.41 0.8764
Bridge 32.93 0.8800 32.86 0.8774 32.93 0.8784 33.05 0.8803 33.15 0.8810 33.05 0.8803 33.14 0.8820 33.27 0.8827
Center 28.77 0.7921 28.71 0.7881 28.79 0.7892 28.88 0.7923 29.00 0.7954 28.88 0.7919 29.03 0.7974 29.09 0.7985
Church 26.30 0.7469 26.41 0.7449 26.46 0.7467 26.52 0.7489 26.59 0.7512 26.56 0.7507 26.64 0.7549 26.70 0.7560
Commercial 29.01 0.7940 29.11 0.7944 29.17 0.7958 29.22 0.7975 29.28 0.7996 29.27 0.7989 29.33 0.8021 29.37 0.8033
D-Residential 24.38 0.6839 24.56 0.6856 24.60 0.6864 24.67 0.6898 24.71 0.6924 24.63 0.6872 24.80 0.6998 24.80 0.6978
Desert 40.20 0.9268 40.22 0.9259 40.17 0.9264 40.06 0.9276 40.31 0.9275 40.24 0.9279 40.27 0.9282 40.25 0.9283
Farmland 35.00 0.8683 34.89 0.8656 34.94 0.8667 34.99 0.8679 35.02 0.8681 34.98 0.8678 35.09 0.8696 35.09 0.8700
Forest 29.85 0.7315 29.88 0.7304 29.90 0.7312 29.99 0.7365 29.98 0.7350 29.97 0.7350 30.05 0.7395 30.05 0.7380
Industrial 28.88 0.7931 28.84 0.7894 28.88 0.7904 28.98 0.7942 29.03 0.7959 28.98 0.7932 29.11 0.7998 29.16 0.8007
Meadow 34.63 0.7804 34.53 0.7769 34.55 0.7784 34.63 0.7814 34.66 0.7807 34.64 0.7813 34.58 0.7807 34.68 0.7820
M-Residential 28.34 0.7365 28.39 0.7347 28.42 0.7349 28.47 0.7370 28.56 0.7401 28.39 0.7334 28.64 0.7443 28.63 0.7436
Mountain 30.63 0.7885 30.70 0.7892 30.70 0.7895 30.74 0.7909 30.78 0.7921 30.75 0.7911 30.79 0.7930 30.79 0.7930
Park 30.54 0.8130 30.63 0.8136 30.65 0.8141 30.72 0.8169 30.76 0.8177 30.73 0.8162 30.81 0.8198 30.82 0.8203
Parking 27.25 0.8317 27.08 0.8245 27.23 0.8270 27.47 0.8352 27.42 0.8354 27.50 0.8363 27.70 0.8435 27.72 0.8424
Playground 35.37 0.8943 35.27 0.892 35.42 0.8929 35.47 0.8942 35.49 0.8946 35.45 0.8946 35.59 0.8968 35.70 0.8976
Pond 32.11 0.8542 32.10 0.8532 32.11 0.8534 32.17 0.8546 32.22 0.8553 32.15 0.8545 32.23 0.8561 32.25 0.8563
Port 28.50 0.8596 28.61 0.8593 28.67 0.8597 28.75 0.8620 28.81 0.8631 28.79 0.8624 28.85 0.8651 28.94 0.8658
Railway Station 28.72 0.7738 28.68 0.7699 28.77 0.7718 28.84 0.7744 28.92 0.7777 28.89 0.7759 28.97 0.7802 29.02 0.7816
Resort 28.52 0.7799 28.59 0.7791 28.62 0.7795 28.68 0.7819 28.74 0.7837 28.66 0.7801 28.78 0.7864 28.82 0.7869
River 31.55 0.7891 31.55 0.7882 31.57 0.7885 31.60 0.7900 31.64 0.7905 31.61 0.7904 31.66 0.7918 31.68 0.7922
School 29.36 0.8044 29.41 0.8033 29.45 0.8041 29.51 0.8067 29.59 0.8091 29.50 0.8048 29.67 0.8123 29.68 0.8124
S-Residential 27.71 0.6728 27.79 0.6723 27.80 0.6725 27.85 0.6752 27.88 0.6758 27.80 0.6728 27.91 0.6782 27.92 0.6775
Square 30.84 0.8200 30.83 0.8181 30.87 0.8183 30.97 0.8218 31.06 0.8236 30.98 0.821 31.11 0.8256 31.15 0.8266
Stadium 29.63 0.8387 29.51 0.834 29.62 0.8358 29.74 0.8388 29.82 0.8413 29.76 0.8394 29.80 0.8420 29.93 0.8439
Storage Tanks 27.44 0.7664 27.50 0.7649 27.52 0.7648 27.58 0.7671 27.63 0.7692 27.58 0.767 27.66 0.7720 27.68 0.7718
Viaduct 28.99 0.7757 28.92 0.7711 28.96 0.7722 29.06 0.7759 29.14 0.7784 29.05 0.7753 29.16 0.7811 29.25 0.7831
Average 30.65 0.8086 30.66 0.8068 30.71 0.8076 30.77 0.8102 30.83 0.8114 30.78 0.8098 30.87 0.8138 30.91 0.8142
Figure 5: Qualitative comparison of state-of-the-art efficient models for ×4 SR task on AID test set.
Figure 6: Qualitative comparison of state-of-the-art efficient SR models for ×4 SR task on DOTA dataset.
Figure 7: Qualitative comparison of state-of-the-art efficient SR models for ×4 SR task on DIOR dataset.
IV Experiments
Datasets. We use the AID dataset [55] as the training set, a large-scale remote sensing benchmark for scene classification, consisting of 30 different scene categories. The AID dataset includes 10,000 HR images, where we randomly select 3,000 for training and 900 for validation. The LR samples are generated by bicubic downsampling. Following TTST [65], we also evaluate our method on the DOTA [56] and DIOR [57] datasets, which contain 900 and 1,000 images, respectively.
Implementation Details. During model training, the learning rate is fixed to $10^{-4}$, and the training procedure stops after 1000 epochs with a batch size of 4. Adam optimizer is used with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. Data augmentation includes random rotations of 90°, 180°, 270°, and horizontal flips on 64×64 patches. The channel number, embedding dimension of cross-attention, and MLP rate of small, medium, and final SpikeSR are set to $\{40,24,72\}$, $\{56,24,72\}$, and $\{64,32,100\}$, respectively. The number of SAGs is 4, with 2 SABs in each SAG. Time step is set to $\rm T=4$.
Metrics. We use the widely used peak signal-to-noise ratio (PSNR) and structural similarity (SSIM) to evaluate SR performance. The results are measured on the Y channel after converting RGB to YCbCr space. For a fair comparison, comparative methods are retained from scratch on the AID dataset, adhering to their official implementation settings.
IV-A Comparison with Efficient Models
We compare our SpikeSR with state-of-the-art (SOTA) efficient SR methods, including CNN-based models of CARN [28], LatticeNet [59], RLFN [61], etc, and transformer-based approaches of SwinIR [27], SPIN [31], HiT-SR [32], etc. We also report results for the small and medium versions of our SpikeSR, denoted as SpikeSR-S and Spike-M, respectively.
Quantitative comparisons. The quantitative results of various methods are reported in Table I. We can observe that our SpikeSR achieves the best performance across three benchmarks, outperforming SOTA CNN- and transformer-based SR models. For example, on AID, DOTA, and DIOR datasets, SpikeSR improves PSNR by 0.08 dB, 0.04 dB, and 0.1 dB, respectively, compared to the impressive SwinIR. Moreover, the small version of SpikeSR requires fewer parameters (472K vs. 897K) and FLOPs (15.21G vs. 23.56G) than SwinIR, yet achieves superior average performance.
Qualitative comparisons. Fig. 5, Fig. 6, and Fig. 7 present visual comparisons on AID, DOTA, and DIOR datasets, respectively. As shown in Fig. 5, SpikeSR effectively restores severely damaged textures, e.g., the runway line in the playground. By contrast, other SR models fail to recover such weak high-frequency details. In Fig. 6, the reconstruction of “img_591” highlights that SpikeSR produces results closest to the GT, while other methods like SPIN recovers unrealistic results. Moreover, Fig. 7 further demonstrates that SpikeSR consistently delivers superior visual quality, restoring more textural information compared to large-capacity ANN-based model like Omni-SR.
V Ablation Study
We conduct ablation studies to assess the impact of key components in SpikeSR. The experimental results in Table III are measured on the AID-tiny dataset [65]. In particular, the Baseline model is constructed by removing the HDA and DSA and replacing them with standard temporal attention (TA), channel attention (CA), and spatial attention (SA) mechanisms. For a fair comparison, we increase the number of $m$, $n$, and channel dimensions to 8, 4, and 256, respectively, which ensures a similar number of parameters with our SpikeSR. Similarly, those settings of Variant-A are modified to 10, 5, and 128, respectively.
TABLE III: Ablation on different variants of our SpikeSR.
Methods TA CA SA HAD DSA #Param. PSNR
Baseline 1120K 27.80
Varian-A 1062K 27.92
Varian-B 1009K 28.11
SpikeSR 1042K 28.19
TABLE IV: Ablation on feature pyramid and deformable convolution.
Methods w/o Pyramid w/o DConv DSA (Full)
#Param. 1042K 918K 1042K
PSNR (dB) 28.14 28.07 28.19
Figure 8: Feature visualizations. The feature obtained by HDA is sharper and preserves more details, indicating high-quality feature representation.
TABLE V: Performance and complexity analysis of SpikeSR with different numbers of blocks.
Number of $m$ 2 3 4 5 6
#Param. 558K 800K 1042K 1284K 1526K
FLOPs 17.47G 25.26G 33.00G 40.84G 48.60G
PSNR (dB) 28.11 28.17 28.19 28.16 28.20
Figure 9: Analysis of local attribution maps (LAMs) [66] and diffusion index (DI). The proposed DAS helps SpikeSR exploit more useful information against CARN and HiT-SR.
Effectiveness of HDA and DSA. Table III indicates how the SR performance is influenced by the HDA and DSA. Comparing the PSNR values of the Baseline and Variant-A reveals that HDA contributes a 0.12 dB improvement. This suggests that enhancing the correlation between the temporal and channel dimensions delivers better recalibration of the membrane potentials. More intuitively, we visualize the feature maps to highlight the impact of HDA, as shown in Fig. 8. The results illustrate that HDA effectively refines the feature representation by emphasizing salient details and suppressing irrelevant information.
By introducing the proposed DSA to the Baseline model, the PSNR results can be improved by a large margin of 0.21 dB. When employing HDA and DSA simultaneously, the resulting SpikeSR achieves an additional 0.08 dB improvement compared to Variant-B. To better demonstrate its effectiveness in capturing global self-similarity priors, we provide the LAM visualization and diffusion index in Fig. 9. As observed, DSA generates more pronounced LAMs and significantly increases the DI, indicating that the model activates more valuable pixels for accurate SR.
Feature Pyramid and DConv. Due to the scale diversity of objects in RSIs, we constructed a feature pyramid to grasp the self-similarity in multiple levels. As listed in Table IV, the use of feature pyramid improves the performance by 0.05 dB without introducing additional parameters. Furthermore, to demonstrate the effectiveness of deformable convolution, we remove this component, which leads to a severe performance drop by 0.12 dB. This illustrates that the self-similar patches contain massive irrelevant and misaligned contents, and direct fusion may introduce interference, thus resulting in suboptimal performance.
Network Depth. We evaluate the impact of the network depth by changing the number of SAGs of our SpikeSR from 2 to 6 blocks. As reported in Table V, SpikeSR achieves the highest SR performance when $m=3$. While increasing the number of $m$ may further improve the reconstruction, it also brings larger model size. Therefore, we set $m=4$ in our final model, considering the trade-off between performance and computational complexity.
VI Conclusion
In this paper, we investigate the application of SNNs for efficient SR of remote sensing images. Motivated by the observation that LIF neurons exhibit a higher spike rate in degraded images, we integrate SNNs with convolutions for improved feature representation. Besides, a hybrid dimension attention is employed to modulate the spike response, further refining salient information. To incorporate valuable prior knowledge for more accurate SR, we propose a deformable similarity attention module, capturing global self-similarity across multiple feature levels. Extensive experiments on various remote sensing datasets demonstrate the efficacy and effectiveness of the proposed SpikeSR model.